### v1

@@ 1 Introduction
Our primary contribution is a novel approach to establishing the existence of a Bayesian-Nash Principals’ Equilibrium (BNPE) in a generalized game with team production and agency concerns. Our methodological approach builds on two complementary strands in the literature on the existence of Bayesian-Nash equilibrium in Bayesian games. [^9] Our work is also related to the literature on the existence of equilibria in discontinuous games initiated by Reny (1999). The related literature is too large to comprehensively summarize here, but some notable contributions include Carmona (2009), Barelli and Meneghel (2013), McLennan, Monteiro, and Tourky (2011), He and Yannelis (2016), Bich and Laraki (2017), Carbonell-Nicolau and McLean (2018), Reny (2020), Olszewski and Siegel (2023), and Prokopovych and Yannelis (2023). The first strand, initiated by Milgrom and Weber (1985), draws on the insight that strategies in Bayesian games can be fully characterized by the joint probability measures they induce over types and actions. Rather than working directly with strategy functions that map types to actions, this distributional approach treats each player’s strategy as a probability measure on the joint type-action space. To determine when two distributional strategies are “close” to each other—and thereby characterize the continuity properties of payoff functions and best-response correspondences—Milgrom and Weber (1985) use the classical narrow topology. [^10] Recall that a sequence of probability measures $\{\mu_{n}\}$ converges to $\mu$ in the narrow topology if and only if $\int fd\mu_{n}\to\int fd\mu$ for every bounded continuous function $f$. When the underlying space is Polish (complete separable metric), this coincides with the weak* topology on the space of probability measures. This topology captures the idea that two probability distributions are close if they assign similar probabilities to similar events, which ensures that a small change in a player’s strategy induces only a small change in all players’ expected payoffs. This notion of closeness is particularly well-suited for Bayesian games, where payoffs depend on the distribution of types and actions.
@@ 3.5 Principals’ Initial Problems
We now turn to the incentive compatibility constraint and the set of incentive compatible mechanisms for team $j$. For notational simplicity, we assume all agents share a common von Neumann-Morgenstern utility function $u$, so that agent heterogeneity enters only through type differences. [^20] All results extend straightforwardly to settings with agent-specific utility functions $u_{i,j}$. We adopt the convention that $u_{i,j}(\widehat{\mathbf{t}},\widehat{\mathbf{a}},\mathbf{w},\mathbf{r}_{j})$ denotes the utility of agent $(i,j)$ when evaluated at type profile $\widehat{\mathbf{t}}$ and action profile $\widehat{\mathbf{a}}$—that is, the subscript $(i,j)$ indicates which agent’s utility is being evaluated. The utility function may depend on agent $(i,j)$’s own type $t_{i,j}$ and action $a_{i,j}$, their team members’ types and actions, and potentially the full profiles $\widehat{\mathbf{t}}$ and $\widehat{\mathbf{a}}$ to capture interdependencies across teams.
@@
Lastly, we specify the payoff functions for the principals. Each team $j$ principal’s von Neumann-Morgenstern utility function is denoted by $\pi_{j}:T^{nN}\times A^{nN}\times W^{N}\to\mathbb{R}$ which may depend on team $j$’s own winnings $w_{j}$, action profile $\mathbf{a}_{j}$, and type profile $\mathbf{t}_{j}$, along with the other teams’ winnings $\mathbf{w}_{-j}$, action profiles $\{\{\mathbf{a}_{j^{\prime}}\}\mid j^{\prime}\neq j\}$ and type profiles $\{\{\mathbf{t}_{j^{\prime}}\}\mid j^{\prime}\neq j\}$.
@@
$\sup_{m_{j}\in IC_{j}(\mathbf{m}_{-j})}E_{\mathbf{m}}\left(\pi_{j}\left(\widehat{\mathbf{t}},\widehat{\mathbf{a}}^{\prime},\mathbf{w}\right)\right)$ (2)
@@ 4 Assumptions and the Metric Structure
The von Neumann-Morgenstern utility function $u$ is a bounded continuous function on $T^{nN}\times A^{nN}\times W^{N}\times I^{n}$ and the von Neumann-Morgenstern utility functions $\{\pi_{j}\}_{j=1}^{N}$ are bounded continuous functions on $T^{nN}\times A^{nN}\times W^{N}$.
@@ 7 Appendix: Proof of Theorem 1
The functional $V_{\pi_{j}}(\mathbf{m})$, defined in equation (8) is continuous in $\mathbf{m}=(m_{j},\mathbf{m}_{-j})$ and concave in $m_{j}=\{\alpha_{j},\kappa_{j}\}$.
@@
Next, the quasi-concavity of $V_{\pi_{j}}$ in $m_{j}$ can be demonstrated as follows. For a given mechanism profile $\mathbf{m}\in M^{N}$, recall the joint action recommendation mechanism component is constructed as:
$\bm{\alpha}^{\mathbf{m}}\!\left(d\widehat{\mathbf{a}}^{\prime}\,\middle|\,\widehat{\mathbf{t}}\right)=\bigotimes_{j=1}^{N}\alpha_{j}\!\left(d\mathbf{a}^{\prime}_{j}\,\middle|\,\mathbf{t}_{j}\right),$
where $\mathbf{m}=\{\alpha_{j},\kappa_{j}\}_{j=1}^{N}$.
Note that the equation (8) expected utility functional for the team $j$ principal may be rewritten as
[…]
For any $m_{j}=(\alpha_{j},\kappa_{j})$, $m_{j}^{\prime}=(\alpha_{j}^{\prime},\kappa_{j}^{\prime})\in M$, and $\lambda\in[0,1]$, define the joint convex combination
$m_{j}^{\lambda}=\big(\alpha_{j}^{\lambda},\kappa_{j}^{\lambda}\big):=\big(\lambda\,\alpha_{j}+(1-\lambda)\,\alpha_{j}^{\prime},\;\lambda\,\kappa_{j}+(1-\lambda)\,\kappa_{j}^{\prime}\big).$
For a given mechanism profile $\mathbf{m}_{-j}\in M^{N-1}$, we construct $\bm{\alpha}^{\mathbf{m}_{-j}}$ as:
$\bm{\alpha}^{\mathbf{m}_{-j}}\!\left(d\widehat{\mathbf{a}}^{\prime}_{-j}\,\middle|\,\widehat{\mathbf{t}}_{-j}\right)=\bigotimes_{j^{\prime}\neq j}\alpha_{j^{\prime}}\!\left(d\mathbf{a}^{\prime}_{j^{\prime}}\,\middle|\,\mathbf{t}_{j^{\prime}}\right).$
Then:
[…]
Since the integral in equation (10) is linear in $\lambda\alpha_{j}+(1-\lambda)\alpha_{j}^{\prime}]$ and $\pi_{j}$ is bounded (Assumption 4), it follows that:
$V_{\pi_{j}}(m_{j}^{\lambda},\mathbf{m}_{-j})=\lambda\,V_{\pi_{j}}(m_{j},\mathbf{m}_{-j})+(1-\lambda)\,V_{\pi_{j}}(m_{j}^{\prime},\mathbf{m}_{-j}),$ (11)
@@
To see this, fix $\mathbf{m}_{-j}$ and suppose $m_{j},m_{j}^{\prime}\in\mathrm{IC}_{j}(\mathbf{m}_{-j})$. For any $\lambda\in[0,1]$, interpret the convex combination $m_{j}^{\prime\prime}=\lambda m_{j}+(1-\lambda)m_{j}^{\prime}$ as a mechanism that implements $m_{j}$ with probability $\lambda$ and $m_{j}^{\prime}$ with probability $1-\lambda$. Then
$f_{i,j}(\mathbf{m}^{\prime\prime})=\lambda f_{i,j}(\mathbf{m})+(1-\lambda)f_{i,j}(\mathbf{m}^{\prime})\geq 0$
whenever both $f_{i,j}(\mathbf{m})\geq 0$ and $f_{i,j}(\mathbf{m}^{\prime})\geq 0$. This holds for all agents $(i,j)$ in team $j$, confirming that $m_{j}^{\prime\prime}\in\mathrm{IC}_{j}(\mathbf{m}_{-j})$.

### v2

@@ 1 Introduction
Our primary contribution is a novel approach to establishing the existence of a Bayesian-Nash Principals’ Equilibrium (BNPE) in a generalized game with team production and agency concerns. Our methodological approach builds on two complementary strands in the literature on the existence of Bayesian-Nash equilibrium in Bayesian games. [^9] Our work is also related to the literature on the existence of equilibria in discontinuous games initiated by Reny (1999). The related literature is too large to comprehensively summarize here, but some notable contributions include Carmona (2009), Barelli and Meneghel (2013), McLennan, Monteiro, and Tourky (2011), He and Yannelis (2016), Bich and Laraki (2017), Carbonell-Nicolau and McLean (2018), Reny (2020), Olszewski and Siegel (2023), and Prokopovych and Yannelis (2023). The first strand, initiated by Milgrom and Weber (1985), draws on the insight that strategies in Bayesian games can be fully characterized by the joint probability measures they induce over types and actions. Rather than working directly with strategy functions that map types to actions, this distributional approach treats each player’s strategy as a probability measure on the joint type-action space. To determine when two distributional strategies are “close” to each other—and thereby characterize the continuity properties of payoff functions and best-response correspondences—Milgrom and Weber (1985) use the classical narrow topology. [^10] Recall that a sequence of probability measures $\{\mu_{n}\}$ converges to $\mu$ in the narrow topology if and only if $\int fd\mu_{n}\to\int fd\mu$ for every bounded continuous function $f$. When the underlying space is Polish (complete separable metric), this coincides with the weak* topology on the space of probability measures. This topology captures the idea that two probability measures are close if they assign similar probabilities to similar events, which ensures that a small change in a player’s strategy induces only a small change in all players’ expected payoffs. This notion of closeness is particularly well-suited for Bayesian games, where payoffs depend on the distribution of types and actions.
@@ 3.5 Principals’ Initial Problems
We now turn to the incentive compatibility constraint and the set of incentive compatible mechanisms for team $j$. For notational simplicity, we assume all agents share a common von Neumann-Morgenstern utility function $u:T^{nN}\times A^{nN}\times W^{N}\times I^{n}\to\mathbb{R}$, so that agent heterogeneity enters only through type differences. [^20] All results extend straightforwardly to settings with agent-specific utility functions $u_{i,j}$. We adopt the convention that $u_{i,j}(\widehat{\mathbf{t}},\widehat{\mathbf{a}},\mathbf{w},\mathbf{r}_{j})$ denotes the utility of agent $(i,j)$ when evaluated at type profile $\widehat{\mathbf{t}}$ and action profile $\widehat{\mathbf{a}}$—that is, the subscript $(i,j)$ indicates which agent’s utility is being evaluated. The utility function may depend on agent $(i,j)$’s own type $t_{i,j}$ and action $a_{i,j}$, their team members’ types and actions, and potentially the full profiles $\widehat{\mathbf{t}}$ and $\widehat{\mathbf{a}}$ to capture interdependencies across teams.
@@
Lastly, we specify the payoff functions for the principals. Each team $j$ principal’s von Neumann-Morgenstern utility function is denoted by $\pi_{j}:T^{nN}\times A^{nN}\times W^{N}\times I^{n}\to\mathbb{R}$ which may depend on team $j$’s own winnings $w_{j}$, action profile $\mathbf{a}_{j}$, type profile $\mathbf{t}_{j}$, and allocation of rewards $\mathbf{r}_{j}$ along with the other teams’ winnings $\mathbf{w}_{-j}$, action profiles $\{\{\mathbf{a}_{j^{\prime}}\}\mid j^{\prime}\neq j\}$ and type profiles $\{\{\mathbf{t}_{j^{\prime}}\}\mid j^{\prime}\neq j\}$.
@@
$\sup_{m_{j}\in IC_{j}(\mathbf{m}_{-j})}E_{\mathbf{m}}\left(\pi_{j}\left(\widehat{\mathbf{t}},\widehat{\mathbf{a}}^{\prime},\mathbf{w},\mathbf{r}_{j}\right)\right)$ (2)
@@ 4 Assumptions and the Metric Structure
The von Neumann-Morgenstern utility function $u$ is a bounded continuous function on $T^{nN}\times A^{nN}\times W^{N}\times I^{n}$ and the von Neumann-Morgenstern utility functions $\{\pi_{j}\}_{j=1}^{N}$ are bounded continuous functions on $T^{nN}\times A^{nN}\times W^{N}\times I^{n}$.
@@ 7 Appendix: Proof of Theorem 1
The functional $V_{\pi_{j}}(\mathbf{m})$, defined in equation (8), is continuous in $\mathbf{m}=(m_{j},\mathbf{m}_{-j})$ and concave in $m_{j}=\{\alpha_{j},\kappa_{j}\}$.
@@
Next, the quasi-concavity of $V_{\pi_{j}}$ in $m_{j}$ can be demonstrated as follows. For any $m_{j}=(\alpha_{j},\kappa_{j})$, $m_{j}^{\prime}=(\alpha_{j}^{\prime},\kappa_{j}^{\prime})\in M$, and $\lambda\in[0,1]$, define the convex combination
$m_{j}^{\lambda}=\lambda m_{j}+(1-\lambda)m_{j}^{\prime}=\big(\lambda\,\alpha_{j}+(1-\lambda)\,\alpha_{j}^{\prime},\;\lambda\,\kappa_{j}+(1-\lambda)\,\kappa_{j}^{\prime}\big).$
Note that this convex combination represents a perfectly correlated randomization over mechanism pairs: with probability $\lambda$ the principal offers mechanism $m_{j}=(\alpha_{j},\kappa_{j})$ and with probability $(1-\lambda)$ offers mechanism $m_{j}^{\prime}=(\alpha_{j}^{\prime},\kappa_{j}^{\prime})$, where the same random draw determines both the action recommendation rule and the reward rule. This ensures that the induced measure on the outcome space is a convex combination of probability measures:
$\mu(m_{j}^{\lambda},\mathbf{m}_{-j})=\lambda\mu(m_{j},\mathbf{m}_{-j})+(1-\lambda)\mu(m_{j}^{\prime},\mathbf{m}_{-j}).$ (9)
Therefore, by linearity of integration with respect to convex combinations of measures:
$V_{\pi_{j}}(m_{j}^{\lambda},\mathbf{m}_{-j})$ $=\int\pi_{j}\,d\mu(m_{j}^{\lambda},\mathbf{m}_{-j})$
$=\int\pi_{j}\,d[\lambda\mu(m_{j},\mathbf{m}_{-j})+(1-\lambda)\mu(m_{j}^{\prime},\mathbf{m}_{-j})]$
$=\lambda\int\pi_{j}\,d\mu(m_{j},\mathbf{m}_{-j})+(1-\lambda)\int\pi_{j}\,d\mu(m_{j}^{\prime},\mathbf{m}_{-j})$
$=\lambda\,V_{\pi_{j}}(m_{j},\mathbf{m}_{-j})+(1-\lambda)\,V_{\pi_{j}}(m_{j}^{\prime},\mathbf{m}_{-j}),$ (10)
@@
To see this, fix $\mathbf{m}_{-j}$ and suppose $m_{j},m_{j}^{\prime}\in\mathrm{IC}_{j}(\mathbf{m}_{-j})$. For any $\lambda\in[0,1]$, the convex combination $m_{j}^{\lambda}=\lambda m_{j}+(1-\lambda)m_{j}^{\prime}$ represents a correlated randomization that offers mechanism $m_{j}$ with probability $\lambda$ and mechanism $m_{j}^{\prime}$ with probability $1-\lambda$. By the same reasoning as in Lemma 1, the induced measure satisfies $\mu(m_{j}^{\lambda},\mathbf{m}_{-j})=\lambda\mu(m_{j},\mathbf{m}_{-j})+(1-\lambda)\mu(m_{j}^{\prime},\mathbf{m}_{-j})$. Therefore, by linearity of the incentive compatibility slack functional:
$f_{i,j}(m_{j}^{\lambda},\mathbf{m}_{-j})=\lambda f_{i,j}(m_{j},\mathbf{m}_{-j})+(1-\lambda)f_{i,j}(m_{j}^{\prime},\mathbf{m}_{-j})\geq 0$
whenever both $f_{i,j}(m_{j},\mathbf{m}_{-j})\geq 0$ and $f_{i,j}(m_{j}^{\prime},\mathbf{m}_{-j})\geq 0$. This holds for all agents $(i,j)$ in team $j$, confirming that $m_{j}^{\lambda}\in\mathrm{IC}_{j}(\mathbf{m}_{-j})$.